OpenStudy (anonymous):

a culture started with 4,000 bacteria. After 5 hours, it grew to 4,800 bacteria. Predict how many bacteria will be present after 15 hours. Round your answer to the nereast whole number. P=Ae^kt

OpenStudy (anonymous):

pleasee help this is the last question i have to pass the test

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

well 5*3=15 so do 800*3 then do 2400+4000= what @saaaaaaammmmmmmmmm

OpenStudy (mathstudent55):

Plug in the known info to find k in the formula. Then rewrite the equation using the value of k you found and solve for P when t = 15. A is the starting amount P is the amount at t time \(\large P = Ae^{kt} \) \(\large 4800 = 4000e^{k \times 5} \) Find k.

OpenStudy (anonymous):

Multiply 400 times 13 hours (400 bacteria every 2 hours), then add the 4000 in addition to it I got 9200 ...you have to round that though, which would be 9000 I believe.